OpenStudy (anonymous):

What kind of triangle is formed if the side lengths are 6, 7, and 8? A. Acute B. Obtuse C. Right D. No triangle can be formed with those side lengths.

OpenStudy (anonymous):

Well we do have a triangle, so D out. They don't follow pythagoras theorem so C out. But we will still use 6^2 + 7^2 = 36+49 = 85. The sum of the square is hence more than 8^2 which is 64. So if this WAS a rt. angled triangle, then the hypotenuse would be root of 85. But 64 is < 85. Hence the triangle formed is acute. A